i need help on this promblem it is 6 and 1half subtract by 1and 9 thenths it is really important it is my homework and i have to turn it in by tomorrow thank you for your time
Answer by checkley71(8403)

6.5-1.9=4.6 IN DECIMAL
OR IN FRACTIONS IT IS
6 1/2 - 1 9/10 CHANGING EACH TO AN IMPROPER FRACTION
13/2 - 19/10 COMMOM DENOMINATOR IS 10
65/10-19/10 = (65-19)/10 OR 46/10 OR 4.6